An open cylindrical tank 4 ft in diameter and 6 ft deep is filled with water and rotated about its axis at 60 rpm. How much liquid is spilled, and how deep is the water at the axis?

1
Expert's answer
2021-11-13T02:09:58-0500


An open cylindrical tank 4 ft in diameter and 6 ft deep is filled with water and rotated about its axis at 60 rpm.

Volume of Tank = πR²h

R = Diameter /2 = 4/2 = 2ft

h = 6ft

= πR²h

= π*2²*6

= 24π ft³

Let say it can be rotated its axis at an angle 60 rpm

Then Volume of water which will not spill = πR²hCos60

= 2πCos60

=-5.98